OpenStudy (anonymous):

3x = 7 – y 2y = 14 – 6x

OpenStudy (anonymous):

both equations can be written as \[y=7-3x\] so they are th e same line

OpenStudy (anonymous):

So y = 7 - 3x put this into second equation 2(7-3x) = 14 - 6x 14 -6x = 14 -6x They are equal and represent the same line on a graph. So infinite number of solutions

OpenStudy (anonymous):

Thanks